# MCGAN: Enhancing GAN Training with Regression-Based Generator Loss

### v1

@@ Abstract
Generative adversarial networks (GANs) have emerged as a powerful tool for generating high-fidelity data. However, the main bottleneck of existing approaches is the lack of supervision on the generator training, which often results in undamped oscillation and unsatisfactory performance. To address this issue, we propose an algorithm called Monte Carlo GAN (MCGAN). This approach, utilizing an innovative generative loss function, termly the regression loss, reformulates the generator training as a regression task and enables the generator training by minimizing the mean squared error between the discriminator’s output of real data and the expected discriminator of fake data. We demonstrate the desirable analytic properties of the regression loss, including discriminability and optimality, and show that our method requires a weaker condition on the discriminator for effective generator training. These properties justify the strength of this approach to improve the training stability while retaining the optimality of GAN by leveraging strong supervision of the regression loss. Numerical results on CIFAR-10 and CIFAR-100 datasets demonstrate that the proposed MCGAN significantly and consistently improves the existing state-of-the-art GAN models in terms of quality, accuracy, training stability, and learned latent space. Furthermore, the proposed algorithm exhibits great flexibility for integrating with a variety of backbone models to generate spatial images, temporal time-series, and spatio-temporal video data.
Keywords Generative adversarial networks  $\cdot$ Image generation $\cdot$ Semi-supervised learning  $\cdot$ Regression
1 Introduction
@@
In recent years, Generative Adversarial Network (GAN) [1] has become one of the most powerful tools for realistic image synthesis. However, the instability of the GAN training and unsatisfying performance remains a challenge. To combat it, much effort has been put into developing regularization methods, see [2, 3, 4, 5]. Additionally, as [6] pointed out, the generator usually suffers gradient vanishing and instability due to the singularity of the denominator showed in the gradient when the discriminator becomes accurate. To address this issue, some work has been done to develop better adversarial loss [7, 8, 9]. As a variant of GAN, conditional GAN (cGAN) [10] is designed to learn the conditional distribution of target variable given conditioning information. It improves the GAN performance by incorporating conditional information to both the discriminator and generator, we hence have better control over the generated samples [11] [12] [13].
@@
Unlike these works on the regularization method and adversarial loss, our work focuses on the generative loss function to enhance the performance of GAN training. In this paper, we propose a novel generative loss, termed as the regression loss $\mathcal{L}_{R}$, which reformulates the generator training as the least-square optimization task on real data. This regression loss underpins our proposed MC-GAN, an enhancement of existing GAN models achieved by replacing the original generative loss with our regression loss. This approach leverages the expected discriminator $D^{\phi}$ under the fake measure induced by the generator. Benefiting from the strong supervision lying in the regression loss, our approach enables the generator to learn the target distribution with a relatively weak discriminator in a more efficient and stable manner.
@@
  • We establish the theoretical foundation of the proposed regression loss, such as the discriminivity and optimality. Besides, we show the improved stability of the proposed MC-GAN. To illustrate, we provide a simple but clear example of Dirac-GAN, in which our proposed MCGAN successfully mitigate the issues of non-convergence of the conventional GANs by incorporating the regression loss.
  • We empirically validate that MCGAN consistently enhances the performance of various GANs across various data, including image data, time series data, and video data. Our approach improves quality, accuracy, training stability, and learned latent space, showing its generality and flexibility.
@@
GANs have demonstrated their capacity to simulate high-fidelity synthetic data, facilitating data sharing and augmentation. Extensive research has focused on designing GAN models for various data types, including images [14], time series [15, 16, 17], and videos [18]. Recently, Conditional GANs (cGANs) have gained significant attention for their ability to generate synthetic data by incorporating auxiliary information [15, 19, 20]. For the integer-valued conditioning variable, conditional GANs can be roughly divided into two groups depending on the way of incorporating the class information: Classification-based and Projection-based cGANs [12, 21, 13, 11, 10, 22]. For the case where conditioning variable is continuous, the training of conditioning GANs is more challenging. For example, conditional WGAN suffers the difficulty in estimating the conditional expected discriminator of real data due to the need for recalibration per every discriminator update [19]. Attempts are made to mitigate this issue, such as conditional SigWGAN [19], which is designed to tackle this issue for time series data.
2 Preliminary
2.1 Generative adversarial networks
@@
Generative adversarial network (GAN) [1] aims to learn the target distribution from real data. To this goal, GAN plays a min-max game between two networks: Generator and Discriminator. Let $\mathcal{X}$ denote the target space and $\mathcal{Z}$ be the latent space. Then Generator, denoted by $G$ is defined as a parameterised function $G$ that maps latent noise $z\in\mathcal{Z}$ to the target data $x\in\mathcal{X}$, i.e. $G:\Theta\times\mathcal{Z}\to\mathcal{X}$. Discriminator denoted by $D$ discriminates between the real data and fake data generated by the generator, it is defined as a parameterised function that maps the data into real space $\mathbb{R}$, i.e., $D:\phi\times\mathcal{X}\to\mathbb{R}$. The generator and discriminator are trained to solve the following min-max game [1]:
$\min_{\theta}\max_{\phi}\mathbb{E}_{\mu}\left[\log(D^{\phi}(X))\right]+\mathbb{E}_{\nu_{\theta}}\left[\log(1-D^{\phi}(X))\right].$ (1)
In Eqn. (1), the discriminator is trained to maximise the binary cross entropy (BCE) loss while the generator is trained to minimise the BCE such that the fake measure induced by generator $\nu_{\theta}$ coincides with $\mu$. Let $\mu$ and $\nu_{\theta}$ denote the true measure and fake measure induced by $G_{\theta}$.
For the convenience of discussion, we can rewrite the GAN training objective (1) in a more generic form as follows:
$\max_{\phi}\mathcal{L}_{D}(\phi;\theta)$ $=\mathbb{E}_{\mu}\left[f_{1}(D^{\phi}(X))\right]+\mathbb{E}_{\nu_{\theta}}\left[f_{2}(D^{\phi}(X))\right],$ (2)
$\min_{\theta}\mathcal{L}_{G}(\theta;\theta)$ $=\mathbb{E}_{\nu_{\theta}}\left[h(D^{\phi}(X))\right],$
@@
where $f_{1}$, $f_{2}$ and $h$ are real-valued functions that can be chosen based on the adversarial loss (or metric) being used. To retrieve the original min-max objective Eqn. (1), we can set $f_{1}(x)=\log(x)$ and $f_{2}(x)=-h(x)=\log(1-x)$. Apart from non-saturating loss, extensive studies are concerned with how to measure the divergence or distance between $\mu$ and $\nu_{\theta}$ as the improved GAN loss function, which are instrumental in stabilising the training and enhancing the generation performance. Examples include Hinge loss[7], Wasserstein loss[9], Least squares loss [8], Energy-based loss [23] among others. Many of them satisfy the general form of Eqn. (2).
Example 1
Here are two examples:
  • Hinge loss : $f_{1}(w)=f_{2}(-w)=-\max(0,1-w)$, and $h(w)=-w$.
  • Wasserstein distance : $f_{1}(w)=f_{2}(-w)=w$, and $h(w)=-w+c_{\mu}$, where $c_{\mu}:=\mathbb{E}_{X\sim\mu}[D^{\phi}(X)]$.
@@
2.2 Conditional GANs
Conditional GAN (cGAN) is a conditional version of a generative adversarial network that can incorporate additional information, such as data labels or other types of auxiliary data into both the generator and discriminative loss [10].
@@ 2.2 Conditional GANs
The goal of conditional GAN is to learn the conditional distribution $\mu$ of the target data distribution $X\in\mathcal{X}$ (i.e., image ) given the conditioning variable (i.e., image class label) $Y\in\mathcal{Y}$. More specifically, under the real measure $\mu$, $X\times Y$ denote the random variable taking values in the space $\mathcal{X}\times\mathcal{Y}$. The marginal law of $X$ and $Y$ are denoted by $P_{X}$ and $P_{Y}$, respectively.
@@
The conditional generator $G_{\theta}:\mathcal{Y}\times\mathcal{Z}\rightarrow\mathcal{X}$ incorporates the additional conditioning variable to the noise input, and outputs the target variable in $\mathcal{X}$. Given the noise distribution $Z$, $G_{\theta}(y)$ induces the induced fake measure denoted by $\nu_{\theta}(y)$, which aims to approximate the conditional law of $\mu(y):=\mathbb{P}(X|Y=y)$ under real measure $\mu$. The discriminator: $D^{\phi}:\mathcal{X}\rightarrow\mathbb{R}$. The task of training an optimal conditional generator is formulated as the following min-max game:
$\mathcal{L}_{D}(\phi,\theta)=\mathbb{E}_{Y\sim P_{Y}}\left[\mathbb{E}_{\mu(Y)}[f_{1}(D^{\phi}(X))]+\mathbb{E}_{\nu_{\theta}(Y)}[f_{2}(D^{\phi}(X)]\right],$ (3)
$\mathcal{L}_{G}(\theta;\phi)=\mathbb{E}_{Y\sim P_{Y}}\left[\mathbb{E}_{X\sim\nu_{\theta}(Y)}[h(D^{\phi}(X))]\right],\enspace\quad\quad\quad\quad\quad\quad\quad$ (4)
@@
where $f_{1}$, $f_{2}$ and $h$ are real value functions as before. Different from the unconditional case, $\mathcal{L}_{D}$ and $\mathcal{L}_{G}$ has in the outer expectation $\mathbb{E}_{y\sim P_{Y}}$ due to $Y$ being a random variable.
3 Monte-Carlo GAN
3.1 Methodology
In this section, we propose the Monte-Carlo GAN (MCGAN) for both unconditional and conditional data generation. Without loss of generality, we describe our methodology in the setting of the conditional GAN task. [^1] The unconditional GAN can be viewed as the conditioning variable is set to be the empty set.
@@ 3.1 Methodology
Consider the general conditional GAN model composed with the generator loss $\mathcal{L}_{G}$ (Eqn. (4)) and the discrimination loss $\mathcal{L}_{D}$ outlined in the last subsection. To further enhance the GAN model, we propose the following MCGAN by replacing the generative loss $\mathcal{L}_{G}$ with the following novel regression loss for training the generator from the perspective of the regression, denoted by $\mathcal{L}_{R}$,
$\mathcal{L}_{R}(\theta;\phi):=\mathbb{E}_{(X,Y)\sim\mu}\left[|D^{\phi}(X)-\mathbb{E}_{x\sim\nu_{\theta}(Y)}[D^{\phi}(x)]|^{2}\right],$ (5)
@@
where the expectation is taken under the joint law $\mu$ of $X$ and $Y$. We optimize the generator’s parameters $\theta$ by minimizing the regression loss $\mathcal{L}_{R}(\theta;\phi)$. We keep the discriminator loss and conduct the remaining min-max training as before.
@@
It is well known that the conditional expectation is the optimal $l^{2}$ estimator. Therefore, the unique minimizer to Eqn (6) is given by the conditional expectation function $f^{*}:\mathcal{Y}\rightarrow\mathbb{R}$, defined as
@@
Minimising the regression loss $\mathcal{L}_{G}$ would enforce the conditional expectation of $D^{\phi}(X)$ under fake measure $\nu_{\theta}(Y)$ to approach that under the conditional true distribution $\mu(Y)=\mathbb{P}(X|Y)$ for any given $D^{\phi}$. Suppose that $(G_{\theta})_{{\theta}\in{\Theta}}$ provides a rich enough family of distributions containing the real distribution $\mu$. Then there exists $\theta^{*}\in\Theta$, which is a minimizer of $\mathcal{L}_{R}(\theta,\phi)$ for all possible discriminator’s parameter $\phi$, and it satisfies that
@@
3.2 Comparison between $\mathcal{L}_{R}$ and $\mathcal{L}_{G}$
@@
$H(\theta,\phi)={\mathbb{E}}_{z\sim p(z)}[\nabla_{\theta}G^{\theta}(z)^{T}\cdot\nabla_{x}D^{\phi}(G^{\theta}(z))].$ (9)
@@
[…]
One can see that Eqn. (10) depends on the discriminator gradients $\nabla_{x}D^{\phi}(G^{\theta_{n}}(z))$ heavily.
MCGAN benefits from the strong supervision of $\mathcal{L}_{R}$, which provides more control over the gradient behaviour during the training. When $\theta$ is close to the optimal $\theta^{*}$, even if $D^{\phi}$ is away from the optimal discriminator, $d_{\phi}(\mu,\nu_{\theta})$ would be small and hence leads to stabilize the generator training. However, it may not be the case for the generator loss as shown in Eq. (10), resulting in the instability of generator training. For example, this issue is evident for the Hinge loss where $h(x)=x$ as shown in [3].
3.3 Illustrative Dirac-GAN example
To illustrate the advantages of MCGAN, we present a toy example from [3], demonstrating its resolution of the training instability in Dirac-GAN. The Dirac-GAN example involves a true data distribution that is a Dirac distribution concentrated at 0. Besides, the Dirac-GAN model consists of a generator with a fake distribution $\nu_{\theta}(x)=\delta(x-\theta)$ with $\delta(\cdot)$ is a Dirac function and a discriminator $D^{\phi}(x)=\phi x$.
@@ 3.3 Illustrative Dirac-GAN example
where $f$ is specified as $f(x)=-\log(1+\exp(x))$. By applying MCGAN to enhance GAN training, the update rules for the MCGAN model parameters $\theta$ and $\phi$ are modified as follows:
@@
Fig. 1 (a-c) demonstrates that GAN, NSGAN and Hinge GAN all fail to converge to obtain the optimal generator parameter $\theta^{*}=0$. That is due to the fact that the updating scheme of $\theta$ depends heavily on the $\phi$. When $\phi$ fails to converge to zero, $\theta$ keeps updating itself even if it has reached zero, and the non-zero $\theta$ further encourages $\phi$ updating away from 00, which results in a vicious cycle and failure of both generator and discriminator. In contrast, Fig. 1(d) of MCGAN training demonstrates that the generator parameter $\theta$ successfully converges to the optimal value 00 thanks to the regression loss in (5) to bring the training stability of the generator.
(a) GAN
(b) NSGAN
(c) HingeGAN
(d) MCGAN
@@
3.4 Discriminability and optimality of MCGAN
To ensure that MCGAN training leads to the optimal generator $\nu_{\theta^{*}}=\mu$, one needs the sufficient discriminative power of $D^{\phi}$. The discriminiative power of the discriminator is determined by the discriminative loss function $\mathcal{L}_{D}$. The discriminative loss function $\mathcal{L}_{D}$ is usually defined as certain divergences, such as JS divergence in GAN [1], and computing such divergence involves finding the optimal discriminator that optimizes the objective function, which might be challenging in practice. See [24] for a comprehensive description of the discriminative loss function.
@@ 3.4 Discriminability and optimality of MCGAN
$\mathcal{P}(\mathcal{X})\times\mathcal{P}(\mathcal{X})\times\mathcal{X}$ $\rightarrow\mathbb{R}$ (11)
$(\mu,\nu,x)$ $\mapsto D^{\phi_{\mu,\nu}}(x),$
@@
The discriminability covers a variety of optimal discriminators in GAN variants. We present in Table 1 a list of optimal discriminators of some commonly used GAN variants along with their values of $a$ and $c$. The detailed description can be found in Appendix A.4.
Although discriminability can be obtained by training the discriminator via certain discriminative loss functions, it is worth emphasizing that the discriminator does not necessarily need to reach its optimum to obtain discriminability. In the case of BCE, the optimal discriminator is given by
However, we can find a non-optimal discriminator with discriminability as follows:
$\hat{D}^{\phi^{*}_{\mu,\nu_{\theta}}}(x)$ $=D^{\phi^{*}_{\mu,\nu_{\theta}}}(x)-\frac{1}{5}\left(D^{\phi^{*}_{\mu,\nu_{\theta}}}(x)-\frac{1}{2}\right)\mathbf{1}_{\{p_{\mu}(x)>p_{\nu_{\theta}}(x)\}}$ (14)
It is easy to verify that $(D^{\hat{\phi}^{*}_{\mu,\nu_{\theta}}}(x)-1/2)(p_{\mu}(x)-p_{\nu_{\theta}}(x))>0,$ and ${D}^{\hat{\phi}^{*}_{\mu,\nu_{\theta}}}(x)\in(0,1)$ for all $x\in\mathcal{X}$ as permitted by the feasible set. Hence, it has discriminability, although it is not optimal. We also present a $1$-D example in Figure 2.
@@
Geometric GAN Hinge loss $12\mathbbm{1}_{\{p_{\mu}(x)\geq p_{\nu_{\theta}}(x)\}}-$ 1 $1$ 00
Energy-based GAN Energy-based loss $m\mathbbm{1}_{\{p_{\mu}(x)<p_{\nu_{\theta}}(x)\}}$ $\operatorname{sign}{(-m)}$ $\frac{m}{2}$
@@
Figure 2: One illustration on the discriminability of the discriminator. In this example, $\mu\sim\mathcal{N}(-2,1),\nu_{\theta}\sim\mathcal{N}(2,1)$. The optimal discriminator is given by (2) and the weaker version of discriminator $\hat{D}^{\phi^{*}}$ is defined in Eqn. (13).
Now we establish the optimality of $\mu=\nu_{\theta}$ in the following theorem under the regularity condition defined in Assumption 3.1.
Assumption 3.1
@@
This assumption is made because that $H(\theta,\phi)=\vec{0}$ outside the equilibrium point usually means the failure of the discriminator (say a constant function), which is of no interest to us. By making this assumption, we focus on the case where the discriminator is powerful enough to guide the generator training.
Theorem 3.1

### v3

@@ Abstract
Generative adversarial networks (GANs) have emerged as a powerful tool for generating high-fidelity data. However, the main bottleneck of existing approaches is the lack of supervision on the generator training, which often results in undamped oscillation and unsatisfactory performance. To address this issue, we propose an algorithm called Monte Carlo GAN (MCGAN). This approach, utilizing an innovative generative loss function, termly the regression loss, reformulates the generator training as a regression task and enables the generator training by minimizing the mean squared error between the discriminator’s output of real data and the expected discriminator of fake data. We demonstrate the desirable analytic properties of the regression loss, including discriminability and optimality, and show that our method requires a weaker condition on the discriminator for effective generator training. These properties justify the strength of this approach to improve the training stability while retaining the optimality of GAN by leveraging strong supervision of the regression loss. Extensive experiments on diverse datasets, including image data (CIFAR-10/100, FFHQ256, ImageNet, and LSUN Bedroom), time series data (VAR and stock data) and video data, are conducted to demonstrate the flexibility and effectiveness of our proposed MCGAN. Numerical results show that the proposed MCGAN is versatile in enhancing a variety of backbone GAN models and achieves consistent and significant improvement in terms of quality, accuracy, training stability, and learned latent space.
Codehttps://github.com/DeepIntoStreams/MCGAN
Introduction
@@
In recent years, Generative Adversarial Network (GAN) (Goodfellow et al. 2014) has become one of the most powerful tools for realistic image synthesis. However, the instability of the GAN training and unsatisfying performance remains a challenge. To combat it, much effort has been put into developing regularization methods, see (Gulrajani et al. 2017; Mescheder, Geiger, and Nowozin 2018; Miyato et al. 2018; Kang, Shin, and Park 2022). Additionally, as (Arjovsky and Bottou 2017) pointed out, the generator usually suffers gradient vanishing and instability due to the singularity of the denominator showed in the gradient when the discriminator becomes accurate. To address this issue, some work has been done to develop better adversarial loss (Lim and Ye 2017; Mao et al. 2017; Arjovsky, Chintala, and Bottou 2017). As a variant of GAN, conditional GAN (cGAN) (Mirza and Osindero 2014) is designed to learn the conditional distribution of target variable given conditioning information. It improves the GAN performance by incorporating conditional information to both the discriminator and generator, we hence have better control over the generated samples (Zhou et al. 2021; Odena, Olah, and Shlens 2017).
@@
Unlike these works on the regularization method and adversarial loss, our work focuses on the generative loss function to enhance the performance of GAN training. In this paper, we propose a novel generative loss, termed as the regression loss $\mathcal{L}_{R}$, which reformulates the generator training as the least-square optimization task. This regression loss underpins our proposed MCGAN, an enhancement of existing GAN models achieved by replacing the original generative loss with our regression loss. This approach leverages the expected discriminator $D^{\phi}$ under the fake measure induced by the generator. Benefiting from the strong supervision lying in the regression loss, our approach enables the generator to learn the target distribution with a relatively weak discriminator in a more efficient and stable manner.
@@
  • We establish the theoretical foundation of the proposed regression loss, e.g., the discriminability, optimality, and improved training stability. A simple but effective toy example of Dirac-GAN is provided to show that our proposed MCGAN successfully mitigates the non-convergence issues of conventional GANs by incorporating regression loss.
  • We empirically validate the consistent improvements of MCGAN over various GANs across diverse data types (i.e., images, time series, and videos). Our approach improves quality, accuracy, training stability, and learned latent space, showing its generality and flexibility.
@@
GANs have demonstrated their capacity to simulate high-fidelity synthetic data, facilitating data sharing and augmentation. Extensive research has focused on designing GAN models for various data types, including images (Han et al. 2018), time series (Yoon, Jarrett, and Van der Schaar 2019; Xu et al. 2020; Ni et al. 2021), and videos (Gupta, Keshari, and Das 2022). Recently, Conditional GANs (cGANs) have gained significant attention for their ability to generate synthetic data by incorporating auxiliary information (Yoon, Jarrett, and Van der Schaar 2019; Liao et al. 2024; Xu et al. 2019). For the integer-valued conditioning variable, conditional GANs can be roughly divided into two groups depending on the way of incorporating the class information: Classification-based and Projection-based cGANs (Odena, Olah, and Shlens 2017; Miyato and Koyama 2018; Kang et al. 2021; Zhou et al. 2021; Mirza and Osindero 2014; Hou et al. 2022). For the case where conditioning variable is continuous, the training of conditioning GANs is more challenging. For example, conditional WGAN suffers difficulty in estimating the conditional expected discriminator of real data due to the need for recalibration per every discriminator update (Liao et al. 2024). Attempts are made to mitigate this issue, such as conditional SigWGAN (Liao et al. 2024), which is designed to tackle this issue for time series data.
Preliminaries
Generative adversarial networks
@@
Generative adversarial networks (GANs) are powerful tools for learning the target distribution from real data to enable the simulation of synthetic data. To this goal, GAN plays a min-max game between two networks: Generator ($G$) and Discriminator ($D$). Let $\mathcal{X}$ denote the target space and $\mathcal{Z}$ be the latent space. Then Generator $G^{\theta}$ is defined as a parameterised function that maps latent noise $z\in\mathcal{Z}$ to the target data $x\in\mathcal{X}$, where $\theta\in\Theta$ is the model parameter of $G$. Discriminator $D^{\phi}:\mathcal{X}\to\mathbb{R}$ discriminates between the real data and fake data generated by the generator.
Let $\mu$ and $\nu_{\theta}$ denote the true measure and fake measure induced by $G^{\theta}$. For generality, the objective functions of GANs can be written in the following general form:
$\max_{\phi}\mathcal{L}_{D}(\phi;\theta)=\mathbb{E}_{\mu}\left[f_{1}(D^{\phi}(X))\right]+\mathbb{E}_{\nu_{\theta}}\left[f_{2}(D^{\phi}(X))\right],$
$\min_{\theta}\mathcal{L}_{G}(\theta;\phi)=\mathbb{E}_{\nu_{\theta}}\left[h(D^{\phi}(X))\right],$ (1)
where $f_{1}$, $f_{2}$ and $h$ are real-valued functions. Different choices of $f_{1}$, $f_{2}$ and $h$ lead to different GAN models.
@@
There are extensive studies concerned with how to measure the divergence or distance between $\mu$ and $\nu_{\theta}$ as the improved GAN loss function, which are instrumental in stabilising the training and enhancing the generation performance. Examples include Hinge loss (Lim and Ye 2017), Wasserstein loss (Arjovsky, Chintala, and Bottou 2017), Least squares loss (Mao et al. 2017), Energy-based loss (Zhao 2016) among others. Many of them satisfy Eqn. (Generative adversarial networks).
Example 1.
  • classical GAN (Goodfellow et al. 2014): $f_{1}(w)=\log(w)$ and $f_{2}(w)=-h(w)=\log(1-w)$.
  • HingeGAN (Lim and Ye 2017): $f_{1}(w)=f_{2}(-w)=-\max(0,1-w)$, and $h(w)=-w$.
  • Wasserstein GAN (Arjovsky, Chintala, and Bottou 2017) : $f_{1}(w)=f_{2}(-w)=w$, and $h(w)=-w+c_{\mu}$, where $c_{\mu}:=\mathbb{E}_{X\sim\mu}[D^{\phi}(X)]$.
@@
Conditional GANs
@@
Conditional GAN (cGAN) is a conditional version of a generative adversarial network that can incorporate additional information, such as data labels or other types of auxiliary data into both the generator and discriminative loss (Mirza and Osindero 2014). The goal of conditional GAN is to learn the conditional distribution $\mu$ of the target data distribution $X\in\mathcal{X}$ (i.e., image ) given the conditioning variable (i.e., image class label) $Y\in\mathcal{Y}$. More specifically, under the real measure $\mu$, $X\times Y$ denote the random variable taking values in the space $\mathcal{X}\times\mathcal{Y}$. The marginal law of $X$ and $Y$ are denoted by $P_{X}$ and $P_{Y}$, respectively.
@@
The conditional generator $G^{\theta}:\mathcal{Y}\times\mathcal{Z}\rightarrow\mathcal{X}$ incorporates the additional conditioning variable to the noise input, and outputs the target variable in $\mathcal{X}$. Given the noise distribution $Z$, $G^{\theta}(y)$ induces the fake measure denoted by $\nu_{\theta}(y)$, which aims to approximate the conditional law of $\mu(y):=P(X|Y=y)$ under real measure $\mu$. The task of training an optimal conditional generator is formulated as the following min-max game:
$\mathcal{L}_{D}(\phi,\theta)=\mathbb{E}_{Y}\left[\mathbb{E}_{\mu(y)}[f_{1}(D^{\phi}(X))]+\mathbb{E}_{\nu_{\theta}(y)}[f_{2}(D^{\phi}(X)]\right],$
$\mathcal{L}_{G}(\theta;\phi)=\mathbb{E}_{Y}\left[\mathbb{E}_{\nu_{\theta}(y)}[h(D^{\phi}(X))]\right],$ (2)
@@
where $f_{1}$, $f_{2}$ and $h$ are real value functions as before and $\mathbb{E}_{Y}$ denotes that the expectation is taken over $y$ sampled from $Y$. Different from the unconditional case, $\mathcal{L}_{D}$ and $\mathcal{L}_{G}$ has in the outer expectation $\mathbb{E}_{y\sim P_{Y}}$ due to $Y$ being a random variable.
Monte-Carlo GAN
Methodology
@@
In this section, we propose the Monte-Carlo GAN (MCGAN) for both unconditional and conditional data generation. Without loss of generality, we describe our methodology in the setting of the conditional GAN task. [^1] The unconditional GAN can be viewed as the conditioning variable is set to be the empty set. Consider the general conditional GAN model composed with the generator loss $\mathcal{L}_{G}$ (Eqn. (Conditional GANs)) and the discrimination loss $\mathcal{L}_{D}$ outlined in the last subsection. To further enhance the GAN model, we propose the following MCGAN by replacing the generative loss $\mathcal{L}_{G}$ with the following novel regression loss for training the generator from the perspective of the regression, denoted by $\mathcal{L}_{R}$,
$\mathcal{L}_{R}(\theta;\phi):=\mathbb{E}_{(x,y)\sim\mu}\left[|D^{\phi}(x)-\mathbb{E}_{\hat{x}\sim\nu_{\theta}(y)}[D^{\phi}(\hat{x})]|^{2}\right],$ (3)
@@
where the expectation is taken under the joint law $\mu$ of $X$ and $Y$. We optimize the generator’s parameters $\theta$ by minimizing the regression loss $\mathcal{L}_{R}(\theta;\phi)$. We keep the discriminator loss and conduct the min-max training as before. The training algorithm of MCGAN is given in Algorithm 1 in Appendix.
@@
It is well known that the conditional expectation is the optimal $l^{2}$ estimator. Therefore, the minimizer to Eqn (4) is given by the conditional expectation function $f^{*}:\mathcal{Y}\rightarrow\mathbb{R}$, defined as
@@
Minimising the regression loss $\mathcal{L}_{G}$ enforces the conditional expectation of $D^{\phi}(X)$ under fake measure $\nu_{\theta}(Y)$ to approach that under the conditional true distribution $\mu(Y)=\mathbb{P}(X|Y)$ for any given $D^{\phi}$. Suppose that $(G^{\theta})_{{\theta}\in{\Theta}}$ provides a rich enough family of distributions containing the real distribution $\mu$. Then there exists $\theta^{*}\in\Theta$, which is a minimizer of $\mathcal{L}_{R}(\theta,\phi)$ for all possible discriminator’s parameter $\phi$, and it satisfies that
@@
Comparison between $\mathcal{L}_{R}$ and $\mathcal{L}_{G}$
@@
$H(\theta,\phi)={\mathbb{E}}_{z\sim P_{Z}}[\nabla_{\theta}G^{\theta}(z)^{T}\cdot\nabla_{x}D^{\phi}(G^{\theta}(z))].$ (7)
@@
$\theta_{n+1}=$ $\theta_{n}-\lambda{\mathbb{E}}_{z\sim P_{Z}}\Big{[}h^{\prime}(D^{\phi}(G^{\theta_{n}}(z)))\nabla_{\theta}G^{\theta}(z)^{T}\Big{|}_{\theta=\theta_{n}}$
$\cdot\nabla_{x}D^{\phi}(G^{\theta_{n}}(z))\Big{]}.$ (8)
One can see that Eqn. (Comparison between $\mathcal{L}_{R}$ and $\mathcal{L}_{G}$) depends on the discriminator gradients $\nabla_{x}D^{\phi}(G^{\theta_{n}}(z))$ heavily.
MCGAN benefits from the strong supervision of $\mathcal{L}_{R}$, which provides more control over the gradient behaviour during the training. When $\theta$ is close to the optimal $\theta^{*}$, even if $D^{\phi}$ is away from the optimal discriminator, $d_{\phi}(\mu,\nu_{\theta})$ would be small and hence leads to stabilize the generator training. However, it may not be the case for the generator loss as shown in Eq. (Comparison between $\mathcal{L}_{R}$ and $\mathcal{L}_{G}$), resulting in the instability of generator training. For example, this issue is evident for the Hinge loss where $h(x)=x$ as shown in (Mescheder, Geiger, and Nowozin 2018).
Illustrative Dirac-GAN example
To illustrate the advantages of MCGAN, we present a toy example from (Mescheder, Geiger, and Nowozin 2018), demonstrating its resolution of the training instability in Dirac-GAN. The Dirac-GAN example involves a true data distribution that is a Dirac distribution concentrated at 0. Besides, the Dirac-GAN model consists of a generator with a fake distribution $\nu_{\theta}(x)=\delta(x-\theta)$ with $\delta(\cdot)$ is a Dirac function and a discriminator $D^{\phi}(x)=\phi x$.
@@
where $f$ is specified as $f(x)=-\log(1+\exp(x))$. By applying MCGAN to enhance GAN training, the update rules for the model parameters $\theta$ and $\phi$ are modified as follows:
@@
Fig. 1 (a-c) demonstrates that GAN, NSGAN and Hinge GAN all fail to converge to obtain the optimal generator parameter $\theta^{*}=0$. That is because the updating scheme of $\theta$ depends heavily on the $\phi$. When $\phi$ fails to converge to zero, $\theta$ continues to update even if it has reached zero, and the non-zero $\theta$ further encourages $\phi$ updating away from 00, which results in a vicious cycle and the failure of both generator and discriminator. In contrast, Fig. 1(d) of MCGAN training demonstrates that the generator parameter $\theta$ successfully converges to the optimal value 00 thanks to the regression loss in (3) to bring the training stability of the generator. A 2D Gaussian mixture example is also provided in Appendix, showing that MCGAN can help mitigate model collapse.
@@
Discriminability and optimality of MCGAN
To ensure that MCGAN training leads to the optimal generator $\nu_{\theta^{*}}=\mu$, one needs the sufficient discriminative power of $D^{\phi}$. The discriminative power of $D^{\phi}$ is determined by the discriminative loss function $\mathcal{L}_{D}$, which is usually defined as certain divergences, such as JS divergence in GAN (Goodfellow et al. 2014). However, computing such divergence involves finding the optimal discriminator that optimizes the objective function, which might be challenging in practice. See (Liu, Bousquet, and Chaudhuri 2017) for a comprehensive description of the discriminative loss function.
@@
$\mathcal{P}(\mathcal{X})\times\mathcal{P}(\mathcal{X})\times\mathcal{X}\rightarrow\mathbb{R};\quad(\mu,\nu,x)\mapsto D^{\phi_{\mu,\nu}}(x),$
@@
Hinge GAN Hinge loss $12\mathds{1}_{\{p_{\mu}(x)\geq p_{\nu_{\theta}}(x)\}}-$ 1 $1$ 00
Energy-based GAN Energy-based loss $m\mathds{1}_{\{p_{\mu}(x)<p_{\nu_{\theta}}(x)\}}$ $\operatorname{sign}{(-m)}$ $\frac{m}{2}$
@@
The discriminability covers a variety of optimal discriminators in GAN variants. We present in Table 1 a list of optimal discriminators of some commonly used GAN variants along with their values of $a$ and $c$. The detailed description can be found in Appendix C. Although discriminability can be obtained by training the discriminator via certain $\mathcal{L}_{D}$, it is worth emphasizing that the discriminator does not necessarily need to reach its optimum to obtain discriminability.
Assumption 1.
@@
Now, we establish the optimality of $\mu=\nu_{\theta}$ in the following theorem under the regularity condition (Assumption 1).
Theorem 1.